OpenStudy (bdflyinyoureye):

(09.04 MC) A quadratic equation is shown below: 9x2 − 16x + 60 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (5 points) Part B: Solve 4x2 + 8x − 5 = 0 by using an appropriate method. Show the steps of your work, and explain why you chose the method used.

KjSaif (kjsaif):

Sorry for your question :(

OpenStudy (gunsodo):

Part A : If complex numbers are not allowed in this problem, the answers do not exist. Because (3x-8/3)^2>=0, always. Which means 9x^2-16x+64/9>=0, which means 9x^2-16x+60 cannot be 0. If complex numbers are allowed, use the formula x = \[\frac{ {-b \pm \sqrt{b^2-4ac}} }{ 2a}\] and you'll be able to determine the complex roots, in the term of i. Part B : This one is pretty much more easier. The equation could be simplify to the form of (2x+5)(2x-1)=0 The reason that I selected this solution is because the coefficients are related to each other. It can be easily simplified to 2 parenthesis.